Suppose an item is manufactured by three machines X, Y and Z. All the machines have equal capacity and are operated at the same rate. It is known that the percentage of defective items produced by X, Y and Z are 2, 7 and 12 percent respectively. All the items produced by X,Y and Z are put into one bin.from this bin one item is drawn at random and is found to be defective. What is the probability that this item was produced by Y?.
1
Expert's answer
2020-06-23T14:48:07-0400

we are already given that the the item picked is defective .

probability = no of favorable outcome /total number of outcomes

And our favorable outcome is that the item is from y, i.e. p(y).

Thus, p(y)= (no of defectives from y)/(total number of defective items)

p(y)= 7/( 2+7+12)

p(y) = 7/21 = 1/3

p(y)= 1/3 = 0.3333